find the other endpoint of the line segment with the given endpoint and midpoint
endpoint: (-7,-1)
midpoint: (6,7)
1
Expert's answer
2019-09-30T05:33:37-0400

X2=2m1-x1=12-(-7)=19

Y2=2m2-y2=14-(-1)=15

Answer: (19,15)